# Does Data Contamination Detection Work (Well) for LLMs? A Survey and Evaluation on Detection Assumptions

### v1

@@ A.5.1 Within-Domain Detection with Different LMs
Table 11: Average contamination detection AUC for the Pythia-12b model, under different PILE domains. ‘PPL’ represents the average perplexity $\pm$ STD, where instances are truncated by the model max length of 2048 tokens.
@@ A.5.2 Metric Distribution in Histogram
We compare between the GitHub and Pile-CC domains. As shown in Figure 2 & 3, when the seen and unseen instances are from the same domain, their PPL_200 distributions are very similar. However, as shown in Figure 4 & 5, when the seen and unseen instances are from different domains, their PPL_200 distributions are very different. This indicates that the PPL_200 relates more to domain shifts, instead of the contamination status of individual instances.
Figure 2: The distribution of PPL_200 when both seen and unseen instances are from the Github domain.
Figure 3: The distribution of PPL_200 when both seen and unseen instances are from the Pile-CC domain.
Figure 4: The distribution of PPL_200 when the seen instances are from the Github domain, and the unseen instances are from the Pile-CC domain.
Figure 5: The distribution of PPL_200 when the seen instances are from the Pile-CC domain, and the unseen instances are from the Github domain.
We observe a similar trend for other metrics: Min 25% Prob (Figure 6, 7, 8, 9), Mem 5 (Figure 10, 11, 12, 13), Entropy 25 (Figure 14, 15, 16, 17).
Figure 6: The distribution of Min 25% Prob when both seen and unseen instances are from the Github domain.
Figure 7: The distribution of Min 25% Prob when both seen and unseen instances are from the Pile-CC domain.
Figure 8: The distribution of Min 25% Prob when the seen instances are from the Github domain, and the unseen instances are from the Pile-CC domain.
Figure 9: The distribution of Min 25% Prob when the seen instances are from the Pile-CC domain, and the unseen instances are from the Github domain.
Figure 10: The distribution of Mem 5 when both seen and unseen instances are from the Github domain.
Figure 11: The distribution of Mem 5 when both seen and unseen instances are from the Pile-CC domain.
Figure 12: The distribution of Mem 5 when the seen instances are from the Github domain, and the unseen instances are from the Pile-CC domain.
Figure 13: The distribution of Mem 5 when the seen instances are from the Pile-CC domain, and the unseen instances are from the Github domain.
Figure 14: The distribution of Entropy 25 when both seen and unseen instances are from the Github domain.
Figure 15: The distribution of Entropy 25 when both seen and unseen instances are from the Pile-CC domain.
Figure 16: The distribution of Entropy 25 when the seen instances are from the Github domain, and the unseen instances are from the Pile-CC domain.
Figure 17: The distribution of Entropy 25 when the seen instances are from the Pile-CC domain, and the unseen instances are from the Github domain.
@@ A.5.3 Cross-Domain Detection with Different Metrics
In this section, we present AUC results with other metrics, when seen and unseen instances are from different domains, for the Pythia-6.9b model. The metrics include Min Top 5% Prob (Figure 18), Mem 1 (Figure 19), and Entropy 5 (Figure 20). All metrics exhibit higher AUC values in the top-right corner and lower values in the bottom-left, while diagonal points approach random guessing.
Figure 18: Average contamination detection AUC for the Pythia-6.9b model with the metric, Min Top 5% Prob, when the seen and unseen instances are from different domains. The abbreviations represent the domains in Table 2.
Figure 19: Average contamination detection AUC for the Pythia-6.9b model with the metric, Mem 1, when the seen and unseen instances are from different domains. The abbreviations represent the domains in Table 2.
Figure 20: Average contamination detection AUC for the Pythia-6.9b model with the metric, Entropy 5, when the seen and unseen instances are from different domains. The abbreviations represent the domains in Table 2.

### v3

@@ A.5.1 Within-Domain Detection with Different LMs
Table 12: Average contamination detection AUC for the pythia-6.9b model, under different domains within the Pile dataset. ‘PPL_200’ represents the average perplexity $\pm$ STD, from the first 200 tokens within every instance. The color green represents AUCs higher than 60.
Assumptions & Metric Github FreeLaw
Enron-
Emails
ArXiv
OpenWeb-
Text2
Open-
Subtitles
Hacker-
News
Youtube-
Subtitles
Pile-CC
A1 PPL_50 49.9 48.7 49.5 51.0 50.4 48.5 48.3 50.7 49.8
PPL_100 50.9 49.6 50.2 50.9 48.9 47.4 47.3 50.2 50.3
PPL_200 50.9 49.5 51.0 51.0 48.5 48.4 50.5 49.4 51.8
Min 5% token 51.9 50.2 49.2 51.4 49.5 49.2 47.8 48.7 50.9
Min 15% token 51.7 49.4 49.8 51.3 49.0 50.2 49.4 48.7 51.0
Min 25% token 51.4 49.1 49.9 51.4 48.6 50.8 49.3 48.8 51.0
A4 Mem 5 49.0 49.0 50.9 53.9 51.0 52.6 49.6 48.5 50.9
Mem 15 48.5 49.5 51.0 52.5 51.1 49.0 49.1 48.3 50.0
Mem 25 49.5 49.2 50.6 52.7 50.7 48.6 49.2 48.2 51.4
A6 Entropy 5 51.0 49.0 49.5 52.0 50.0 49.2 48.8 49.3 51.0
Entropy 15 50.9 49.1 49.7 51.9 49.6 50.6 48.8 49.1 50.9
Entropy 25 50.9 49.2 49.8 51.8 49.4 51.0 48.8 49.0 50.9
Average AUC 50.3 49.2 50.2 52.1 49.9 49.9 48.9 48.9 50.8
PPL_200 Seen 2.6$\pm$2.8 3.4$\pm$1.7 5.4$\pm$4.0 7.5$\pm$3.5 10.4$\pm$5.4 11.0$\pm$5.7 11.2$\pm$4.0 12.3$\pm$15.0 12.9$\pm$8.4
Unseen 2.7$\pm$3.7 3.6$\pm$5.6 5.9$\pm$5.0 7.5$\pm$3.4 10.1$\pm$5.0 10.6$\pm$5.0 11.5$\pm$5.0 11.8$\pm$10.0 13.4$\pm$10.1
Table 13: Average contamination detection AUC for the pythia-12b model, under different domains within the Pile dataset. ‘PPL_200’ represents the average perplexity $\pm$ STD, from the first 200 tokens within every instance. The color green represents AUCs higher than 60.
Assumptions & Metric cpp python Github-Lean julia tex Github-Isabelle fortran Github-Coq r
A1 PPL_50 51.1 51.7 49.9 50.2 49.4 50.9 50.3 49.1 49.5
PPL_100 51.4 51.6 52.0 51.3 50.4 51.7 50.6 49.1 51.4
PPL_200 51.8 50.8 51.2 51.6 50.0 50.8 51.1 49.0 53.0
Min 5% token 50.1 51.1 49.0 52.0 49.4 51.2 48.1 49.4 50.9
Min 15% token 50.2 50.4 50.7 51.2 48.6 50.4 48.1 49.0 51.5
Min 25% token 50.3 49.8 51.3 51.0 48.7 49.9 48.2 48.9 52.4
A4 Mem 5 51.3 48.2 51.8 49.0 48.2 55.4 50.8 50.5 51.6
Mem 15 52.0 49.6 50.7 48.3 48.5 54.6 50.6 50.4 52.0
Mem 25 51.6 50.5 50.1 49.1 49.4 54.3 50.4 50.8 50.9
A6 Entropy 5 50.2 49.4 51.7 51.3 48.8 49.0 48.8 48.9 54.1
Entropy 15 50.2 49.4 51.7 51.3 48.8 48.9 48.8 49.0 53.5
Entropy 25 50.3 49.5 51.8 51.2 48.8 49.0 48.8 49.0 53.5
Average AUC 50.9 49.8 50.9 50.4 49.0 51.6 49.6 49.7 52.2
PPL_200 Seen 4.3$\pm$2.6 6.7$\pm$3.9 6.9$\pm$3.4 8.3$\pm$5.0 8.3$\pm$5.1 8.7$\pm$5.4 9.5$\pm$6.8 10.4$\pm$8.3 10.5$\pm$7.0
Unseen 4.6$\pm$3.0 6.8$\pm$4.1 6.9$\pm$3.0 8.6$\pm$5.6 8.7$\pm$6.3 9.2$\pm$6.5 9.9$\pm$7.6 9.9$\pm$7.2 10.5$\pm$5.6
Table 14: Average contamination detection AUC for the OLMo-2-1124-7B model, under different domains within the Algebraic Stack dataset. ‘PPL_200’ represents the average perplexity $\pm$ STD, from the first 200 tokens within every instance. The color green represents AUCs higher than 60.
Assumptions & Metric
RE-
2012temp
STS-
B
DC-
MTSample
RE-
2011coref
events-
BioRed
events-
NLMGene
events-
2012temp
events-
2006deid
events-
BioASQ
A1 PPL_50 62.5 83.3 51.9 60.5 50.4 49.4 50.0 51.4 67.6
PPL_100 95.3 93.3 60.0 70.0 50.0 48.8 50.4 59.3 52.6
PPL_200 99.4 96.8 58.9 70.5 89.4 87.4 76.2 79.0 66.1
Min 5% token 92.9 93.4 47.4 72.6 61.4 76.1 57.5 78.0 74.9
Min 15% token 93.3 93.4 51.7 70.1 88.3 85.3 71.2 82.3 69.4
Min 25% token 93.4 92.1 53.3 68.5 94.1 80.5 74.8 76.5 64.7
A4 Mem 5 41.4 48.0 47.9 46.6 49.8 52.0 51.0 52.5 49.2
Mem 15 45.2 53.3 49.1 46.4 48.9 52.3 51.9 52.7 52.8
Mem 25 48.9 55.9 50.3 48.4 46.6 52.1 52.6 53.1 52.5
A6 Entropy 5 93.2 80.5 55.4 63.9 94.5 65.3 74.3 62.6 43.1
Entropy 15 93.1 82.0 54.8 64.5 94.5 66.6 74.2 63.7 46.0
Entropy 25 93.1 82.6 54.1 64.7 94.5 67.1 74.4 64.2 47.6
Average AUC 74.1 73.1 52.0 59.0 69.8 63.5 62.4 62.8 55.1
PPL_200 Seen 1.4$\pm$0.1 1.5$\pm$0.1 1.7$\pm$0.2 2.1$\pm$0.8 2.8$\pm$0.3 3.1$\pm$0.4 3.2$\pm$0.4 3.3$\pm$0.6 8.1$\pm$2.3
Unseen 3.0$\pm$1.6 2.0$\pm$0.3 1.8$\pm$0.2 3.6$\pm$2.4 3.8$\pm$0.9 4.3$\pm$1.0 4.1$\pm$1.1 4.6$\pm$1.6 9.4$\pm$2.7
Table 15: Average contamination detection AUC for the BioMistral model, under different domains within the Medical-NLU dataset. ‘PPL_200’ represents the average perplexity $\pm$ STD, from the first 200 tokens within every instance. The color green represents AUCs higher than 60.
@@ A.5.2 Metric Distribution in Histogram
We compare the MIA performance between the GitHub and Pile-CC domains, from the Pythia-6.9b model. As shown in Figure 2 & 3, when the seen and unseen instances are from the same domain, their PPL_200 distributions are very similar. However, as shown in Figure 4 & 5, when the seen and unseen instances are from different domains, their PPL_200 distributions are very different. This indicates that the PPL_200 relates more to domain shifts, instead of the contamination status of individual instances.
Figure 2: The density plot of PPL_200 from the Pythia-6.9b model, when both seen and unseen instances are from the Github domain.
Figure 3: The density plot of PPL_200 from the Pythia-6.9b model, when both seen and unseen instances are from the Pile-CC domain.
Figure 4: The density plot of PPL_200 from the Pythia-9.6b model, when the seen instances are from the Github domain, and the unseen instances are from the Pile-CC domain.
Figure 5: The density plot of PPL_200 from the Pythia-9.6b model, when the seen instances are from the Pile-CC domain, and the unseen instances are from the Github domain.
We observe a similar trend for other metrics: Min 25% Prob (Figure 6, 7, 8, 9), Mem 25 (Figure 10, 11, 12, 13), Entropy 25 (Figure 14, 15, 16, 17).
Figure 6: The density plot of Min 25% Prob when both seen and unseen instances are from the Github domain.
Figure 7: The density plot of Min 25% Prob when both seen and unseen instances are from the Pile-CC domain.
Figure 8: The density plot of Min 25% Prob from the Pythia-9.6b model, when the seen instances are from the Github domain, and the unseen instances are from the Pile-CC domain.
Figure 9: The density plot of Min 25% Prob from the Pythia-9.6b model, when the seen instances are from the Pile-CC domain, and the unseen instances are from the Github domain.
Figure 10: The density plot of Mem 25 from the Pythia-6.9b model, when both seen and unseen instances are from the Github domain.
Figure 11: The density plot of Mem 25 from the Pythia-6.9b model, when both seen and unseen instances are from the Pile-CC domain.
Figure 12: The density plot of Mem 25 from the Pythia-9.6b model, when the seen instances are from the Github domain, and the unseen instances are from the Pile-CC domain.
Figure 13: The density plot of Mem 25 from the Pythia-9.6b model, when the seen instances are from the Pile-CC domain, and the unseen instances are from the Github domain.
Figure 14: The density plot of Entropy 25 from the Pythia-6.9b model, when both seen and unseen instances are from the Github domain.
Figure 15: The density plot of Entropy 25 from the Pythia-6.9b model, when both seen and unseen instances are from the Pile-CC domain.
Figure 16: The density plot of Entropy 25 from the Pythia-9.6b model, when the seen instances are from the Github domain, and the unseen instances are from the Pile-CC domain.
Figure 17: The density plot of Entropy 25 from the Pythia-9.6b model, when the seen instances are from the Pile-CC domain, and the unseen instances are from the Github domain.
@@ A.5.3 Cross-Domain Detection with Different Metrics
In this section, we present AUC results with other metrics, when seen and unseen instances are from different domains, for the Pythia-6.9b model. The metrics include Min 25% token (Figure 18), Mem 25 (Figure 19), and Entropy 25 (Figure 20). All metrics exhibit higher AUC values in the top-right corner and lower values in the bottom-left, while diagonal points approach random guessing.
Figure 18: Average contamination detection AUC for the Pythia-6.9b model with the metric, Min 25% token, when the seen and unseen instances are from different domains. The abbreviations represent the domains in Table 12.
Figure 19: Average contamination detection AUC for the Pythia-6.9b model with the metric, Mem 25, when the seen and unseen instances are from different domains. The abbreviations represent the domains in Table 12.
Figure 20: Average contamination detection AUC for the Pythia-6.9b model with the metric, Entropy 25, when the seen and unseen instances are from different domains. The abbreviations represent the domains in Table 12.